### v1

# Retrofitting Small Multilingual Models for Retrieval: Matching 7B Performance with 300M Parameters

Abstract
Training effective multilingual embedding models presents unique challenges due to the diversity of languages and task objectives. Although small multilingual models (<1 B parameters) perform well on multilingual tasks generally, they consistently lag behind larger models (>1 B) in the most prevalent use case: retrieval. This raises a critical question: Can smaller models be retrofitted specifically for retrieval tasks to enhance their performance? In this work, we investigate key factors that influence the effectiveness of multilingual embeddings, focusing on training data scale, negative sampling strategies, and data diversity. We find that while increasing the scale of training data yields initial performance gains, these improvements quickly plateau—indicating diminishing returns. Incorporating hard negatives proves essential for consistently improving retrieval accuracy. Furthermore, our analysis reveals that task diversity in the training data contributes more significantly to performance than language diversity alone. As a result, we develop a compact (approximately 300M) multilingual model that achieves retrieval performance comparable to—or even surpassing—current strong 7B models. [^1] This draft was completed and patented in May 2025.
Retrofitting Small Multilingual Models for Retrieval: Matching 7B Performance with 300M Parameters
Lifu Tu, Yingbo Zhou, Semih Yavuz Salesforce AI Research
1 Introduction
Text embedding models (Devlin et al., 2019; Xue et al., 2021; Izacard et al., 2022) provide dense vector representations that capture semantic meaning, enabling efficient retrieval of relevant passages or documents from large corpora based on a given query. The importance of high-quality embeddings has grown significantly with the rise of retrieval-augmented generation (RAG) systems (Gao et al., 2024), where retrieval performance directly impacts overall system effectiveness.
Recent benchmarks, such as MMTEB (Enevoldsen et al., 2025)—a comprehensive test suite covering over 100 embedding evaluation tasks across more than 250 languages—demonstrate that compact multilingual models can perform well across a range of tasks. For instance, multilingual-e5-large-instruct, with just 560 million parameters, is the best-performing publicly available model on MMTEB overall. However, in the most common use case—retrieval—the best small models still lag significantly behind large language models (LLMs) with several billions of parameters, typically underperforming by around 2 points in retrieval metrics.
In this work, we aim to significantly enhance the retrieval performance of small-scale multilingual embedding models—those with around 300 million parameters—bringing them on par with, or even surpassing, the performance of current strong 7B models. To accomplish this, we investigate a strategy for retrofitting these compact models using synthetic multilingual training data derived from the mC4 corpus (Xue et al., 2021). Our approach involves systematically exploring several key factors that influence the effectiveness of multilingual embedding training, including the choice of data sources, the scale of training data, and negative sampling strategies. Through extensive experiments, we uncover several important findings. First, we demonstrate that synthetic multilingual data—when carefully curated—can substantially improve retrieval performance, validating its value as a scalable training resource. Second, we observe that increasing the amount of training data yields substantial performance gains initially, but these gains diminish as the data scale continues to grow, suggesting a point of saturation. Third, we find that mining hard negatives using a strong backbone embedding model consistently enhances retrieval accuracy across languages. Lastly, our analysis highlights the importance of both task diversity and language diversity in training data, with task diversity showing a particularly strong impact on downstream performance.
Finally, our multilingual embedding model (approximately 300M parameters) achieves a score of 60.56 on the MMTEB (Multilingual) retrieval task category—outperforming or matching the performance of current strong 7B models. Notably, our model also performs competitively across other task categories, achieving a mean score that is nearly 1 point higher than comparable models on average across all MMTEB tasks.
2 Related work
Multilingual Embedding Models.
Text embeddings are fundamental to many downstream NLP tasks, such as semantic similarity and information retrieval. Their importance has only increased with the advent of RAG systems, where embedding quality directly affects overall performance. Recently, several studies have leveraged large language models (LLMs) as embedding models (Wang et al., 2023; Meng et al., 2024; Yu et al., 2024; Lee et al., 2025), initializing from pre-trained architectures like XLM-R (Conneau et al., 2020), and Mistral-7B (Jiang et al., 2023). These works demonstrate clear benefits of LLM-based initialization.Although smaller models—such as multilingual-e5-large-instruct have emerged as the best publicly available options on benchmarks like MMTEB (Enevoldsen et al., 2025), their retrieval performance still trails significantly behind that of larger 7B+ models. This gap limits the adoption of compact models in retrieval-centric applications, including RAG systems. In this work, we address this limitation by developing a highly effective small multilingual embedding model that rivals or exceeds the performance of current 7B models.
Synthetic Data Generation.
Recent work (Xiao et al., 2023; Lee et al., 2025) has explored using large language models (LLMs) to generate synthetic data from documents or passages, aiming to enhance the capabilities of multilingual embedding models. In some cases, additional filtering techniques (Thakur et al., 2024; Lee et al., 2025) are employed to improve the quality of the generated data. In our work, we utilize clean Wikipedia documents from mC4 (Xue et al., 2021) as source material and employ a strong generation model (GPT-4o-mini) to create synthetic query-document pairs. Our analysis confirms that the sampled synthetic data is of high quality and contributes meaningfully to model performance.
3 Method
Data.
We use the existing high-quality diverse English data from different task category: MS MARCO (Nguyen et al., 2016), HotpotQA (Yang et al., 2018), NQ (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), AllNLI from SimCSE (Gao et al., 2021) and other task type data can be founds in Appendix D.
We also generate synthetic data to mitigate the shortage of multilingual retrieval data. Specifically, we sample lengthy (100-1000) Wikipedia articles from mC4222https://huggingface.co/datasets/allenai/c4, then genera a question given the Wikipedia documents with GPT-4o-mini. The generated question and the sampled article constitute a text pair to the fine-tuning data. We generate synthetic data across 20 languages [^3] Please see Appendix A.1 for the full language list.
Training.
Given a batch of query document pair, we use the contrastive loss (Gao et al., 2021):
$\mathrm{L}=-\frac{1}{N}\Bigg($
$\sum_{i=1}^{N}\log\frac{\exp(\frac{\textit{sim}(h_{q_{i}},h_{d^{+}})}{\tau)}}{\exp(\frac{\textit{sim}(h_{q_{i}},h_{d^{+}})}{\tau})+\sum_{j=1}^{K}\exp(\frac{\textit{sim}(h_{q_{i}},h_{d_{j}^{-}})}{\tau})}\Bigg)$
where $q_{i}$ is the $i$-th query of the mini batch, $d^{+}$ is the positive document, and $d_{j}^{-}$ is the negative document. $\tau$ is the temperature term, $N$ is the mini batch size, $K$ is the number of negative documents associated with each query. $h_{q}$ and $h_{d}$ are the representations of the given query $q$ and document $d$, sim is the similarity function. Cosine similarity is used in our experiments. We set $\tau=0.02$ in our experiments. More experimental details are shown in the Appendix C.
Evaluation.
We mainly use MMTEB (Enevoldsen et al., 2025) for the multilingual evaluation , which consists of a large collection of more than 500 distinct tasks covering 250+ languages and multiple domains.
4 Results
Model #Params Supported Context Length Emb. Dim. MMTEB(Retrieval)
MGTE 305M 8k 768 56.50
bge-m3 568M 8k 4096 55.59
multilingual-e5-large-instruct 560M 514 1024 57.12
snowflake-l 568M 8k 1024 58.36
Our 305M 8k 768 60.56
gte-Qwen2-1.5B-instruct 1.5B 32k 1536 60.78
SFR-Embedding-Mistral 7B 4k 4096 59.44
gte-Qwen2-7B-instruct 7B 32k 3584 60.08
Table 1: Performance of models in the MTEB (Multilingual) retrieval tasks.
As shown in prior work (Enevoldsen et al., 2025), multilingual-e5-large-instruct is the best-performing publicly available model on the MTEB (Multilingual) benchmark. However, it only supports a context length of approximately 500 tokens. Despite its strong performance, its retrieval score is over 3 points lower than that of gte-Qwen2-1.5B-instruct. Furthermore, as shown in Table 1, on MTEB (Multilingual) retrieval tasks, the best-performing small multilingual model still lags behind larger models by about 2 points. In this work, we present a compact embedding model (about 300M parameters) that achieves the best performance among small multilingual models on retrieval benchmarks—comparable to or even exceeding that of significantly larger models.
5 Analysis
5.1 Data Resource
In this section, we first examine the usefulness of multilingual synthetic data for retrieval. To explore this, we sample 2,000 examples from three different sources: high-quality English retrieval data from MS MARCO (Nguyen et al., 2016), English–French parallel data from OPUS (Tiedemann, 2012), and our own sampled French synthetic query-document pairs. We evaluate using two backbone embedding models: a multilingual model (MGTE (Zhang et al., 2024)) and an English-only model (BGE-en-1.5 (Xiao et al., 2023)). As shown in Table 2, our synthetic data significantly improves multilingual retrieval performance. On the other hand, parallel data contributes little to retrieval effectiveness.
Setting all English Multilingual
MGTE 56.50 49.80 67.44
+English Data 58.6 52.51 68.21
+parallel data 57.24 50.68 67.54
+synthetic data 58.83 52.6 68.81
(a) Multilingual Embedding Model
Setting all English Multilingual
BGE-en-1.5 38.02 45.73 25.91
+English Data 39.30 46.26 26.87
+parallel data 37.67 45.76 23.73
+synthetic data 42.53 49.26 31.36
(b) Non-Multilingual Embedding Model
Table 2: Performance on MTEB (Multilingual) when fine-tuning with different data sources. The column ‘all’, ‘English’, and ‘Multilingual’ refer to the average performance across all retrieval tasks, English-only tasks, and tasks involving non-English languages, respectively.
5.2 Data Scale Effect
In the previous subsection, we observed that our synthetic data significantly improves multilingual retrieval performance. This raises an important question: Is it efficient to continue improving retrieval performance by simply adding more synthetic data? To explore this, we sampled French synthetic retrieval data at varying sizes: 1k, 2k, 4k, 8k, and 16k examples. As shown in Figure 1, multilingual retrieval performance continues to improve with increased data size. However, the performance gains diminish after 4k examples, with each subsequent data doubling yielding smaller improvements compared to the earlier stages. Based on this observation, we choose to generate 4k synthetic samples per language in the following experiments.
Figure 1: Synthetic Data scale experiments.
5.3 Hard Negatives
Recent studies (Lee et al., 2024, 2025) have demonstrated the effectiveness of incorporating hard negatives in training embedding models. In this section, we compare two strategies for selecting hard negatives: (a) vanilla: neegatives are randomly sampled within each mini-batch. (b) hard negatives: negatives are mined using the backbone multilingual embedding model.
As shown in Figure 2, the hard negative strategy consistently improves multilingual retrieval performance. Based on this observation, we adopt the hard negative mining approach for generating synthetic training samples in our experiments.
Figure 2: Hard negatives experiments.
5.4 Task Diversity and Language Diversity
Mul. Eur. Fr Ja
Our (all) 60.56 59.76 68.28 72.29
En-Mix 59.60 58.22 67.56 72.26
Mul-Syn 59.36 58.61 67.19 71.46
En-Syn + En-Mix 60.38 59.56 68.05 72.29
Table 3: Average retrieval performance across different levels of task and language diversity. "Mul.", "Eur.", "Fr", and "Ja" refer to the average scores on MTEB (Multilingual) retrieval tasks, MTEB (Europe) retrieval tasks, 9 French retrieval tasks, and 10 Japanese retrieval tasks, respectively.
As shown in Table 3, our multilingual embedding model achieves a score of 59.36 on the MTEB (Multilingual) retrieval tasks when trained with all synthetic data (Mul-Syn)—a significant improvement over the backbone model baseline (56.5). To further enhance performance, we train a model using all English data (En-Mix) across a variety of task types (e.g., classification, STS, clustering). This model performs strongly, highlighting the value of diverse task data. By combining all synthetic multilingual data with the English diverse task data, our final model reaches a score of 60.56, improving by another full point.
Interestingly, when training only with English synthetic data and English diverse task data (i.e., no multilingual synthetic data) (Syn-En + En-Mix), the resulting model scores just 0.2 points lower on the MTEB (Multilingual) retrieval tasks. To better understand this performance difference, we also report results on MTEB (Europe), as well as on 9 French and 10 Japanese retrieval tasks (listed in the Appendix E). These results show minimal variation across the two models on high-resource languages, suggesting that task diversity contributes more to performance gains than language diversity when continuing to retrofit multilingual embedding models.
5.5 Performance Across Other Task Categories
Table 4 shows that our models achieve over 1 point higher overall performance and outperform the baseline across most task types. This improvement suggest the effectiveness of our training strategy—particularly the incorporation of diverse task types and synthetic multilingual data—in producing more robust and generalizable multilingual embeddings. The consistent gains across multiple task categories suggest that the model is not overfitting to any single task or language, but is instead learning semantically rich representations that transfer well to a variety of scenarios.
Our MGTE
Mean (Type) 52.66 51.51
- Bitext Mining 71.61 71.79
- Classification 58.98 57.17
- Clustering 46.11 44.96
- Inst. Retrieval -0.85 -0.74
- Multilabel Class. 20.84 19.82
- Pair Class. 81.01 80.49
- Reranking 61.70 60.72
- Retrieval 60.56 56.50
- STS 73.96 72.85
Table 4: Comparison with the backbone model (MGTE) on the MTEB (Multilingual) benchmark for massive multilingual embedding.
6 Conclusion
In this work, we develop a compact multilingual embedding model (approximately 300M parameters) that achieves retrieval performance comparable to—or even surpassing—that of current strong 7B models. This demonstrates that small models can be effectively used for retrieval, the most critical and practical application of embedding models.
Limitations
In this work, we explore methods to improve the retrieval performance of small multilingual models and demonstrate consistent gains across most task types. However, several limitations remain:
Benchmark
MTEB (Multilingual) covers over 250 languages; however, low-resource languages are primarily represented in task categories such as bitext mining and classification. The evaluation of retrieval performance for these languages remains limited, which constrains our ability to assess the effectiveness of the proposed methods in low-resource settings. This limitation may also explain why the improvement from multilingual synthetic data—compared to English-only synthetic data—is relatively small, as shown in Table 3.
Efficiency
In this work, we observe that multilingual synthetic data is effective in improving retrieval performance. However, its benefits eventually plateau, revealing a bottleneck when relying solely on such data. We find that further improvements can be achieved by incorporating diverse non-retrieval task types. Exploring how to efficiently leverage both non-retrieval task data and multilingual synthetic query-document pairs presents a promising direction for future research.
Appendix A Details on Synthetic Data Generation
A.1 Languages for Synthetic Data
List of language ISO-639 code (https://cloud.google.com/translate/docs/languages): en, fr, es de,zh, it, ja, ko, fa, hi, id, ar, bn, fi, sw, te, th, jv, ms, sq.
A.2 Template for Synthetic Data Generation
. Give a sampled Wikipedia documents {d} and target language {lang}, the following template is used to generate a query:
You are a curious AI assistant, please generate one specific and valuable question based on the following text. If the text is not {lang}, please reply with Non-{lang}. The generated question should revolve around the core content of this text, and avoid using pronouns (e.g., ’this’). Note that you should generate only one question, without including additional content:
{d}
Appendix B Models
Below is a list of models discussed in this work:
  • MGTE (Zhang et al., 2024) [^4] Alibaba-NLP/gte-multilingual-base
  • BGE-en-1.5 (Xiao et al., 2023) [^5] BAAI/bge-large-en-v1.5
  • bge-m3 (Chen et al., 2024) [^6] BAAI/bge-m3
  • multilingual-e5-large-instruct (Wang et al., 2024)777intfloat/multilingual-e5-large-instruct
  • snowflake-l (Yu et al., 2024) [^8] Snowflake/snowflake-arctic-embed-l-v2.0
  • gte-Qwen2-1.5B-instruct (Li et al., 2023) [^9] Alibaba-NLP/gte-Qwen2-1.5B-instruct
  • SFR-Embedding-Mistral (Meng et al., 2024) [^10] Salesforce/SFR-Embedding-Mistral
  • gte-Qwen2-7B-instruct (Li et al., 2023) [^11] Alibaba-NLP/gte-Qwen2-7B-instruct
Appendix C Implementation Details
We use a batch size of 1024 and a learning rate of 1e-5, with 100 warm-up steps followed by linear decay. The maximum sequence lengths are set to 256 for queries and 512 for documents. LoRA adapters with a rank of $r$ = 8 are applied. Each query-document pair is trained with 7 hard negatives, all drawn from the same task.
Appendix D Additional Data Details
The following are additional English data samples from the training split of MTEB, excluding data used in the official MTEB evaluation.
  • Classification: AmazonPolarity, AmazonReviews, Imdb, MTOPDomainClassification,AmazonReviewsClassification
  • STS: STS12, STS22, STSBenchmark
  • Reranking: SciDocsRR and StackOverflowDupQuestions
  • Clustering: arXIV, bioRxiv, medRxiv
  • Pair Classification: SprintDuplicateQuestions, TwitterURLCorpus
Appendix E Additional Retrieval Tasks
9 French retrieval tasks:
French subset of the following 9 tasks: AlloprofRetrieval, BSARDRetrieval, FQuADRetrieval, SyntecRetrieval, BelebeleRetrieval, MIRACLRetrievalHardNegatives, PublicHealthQA, StatcanDialogueDatasetRetrieval, WebFAQRetrieval
10 Japanese retrieval tasks:
Japanese subset of the following 10 tasks: JaGovFaqsRetrieval, JaqketRetrieval, JaQuADRetrieval, NLPJournalAbsIntroRetrieval, NLPJournalTitleAbsRetrieval, NLPJournalTitleIntroRetrieval, BelebeleRetrieval, MIRACLRetrievalHardNegatives, MrTidyRetrieval, WebFAQRetrieval.