### v2

# Retrofitting Small Multilingual Models for Retrieval: Matching 7B Performance with 300M Parameters

Abstract
Training effective multilingual embedding models is challenging due to linguistic diversity and varying task objectives. While small multilingual models (fewer than 1B parameters) perform well on general multilingual benchmarks, they lag behind larger models in retrieval tasks—the dominant real-world application of embeddings. In this work, we investigate whether smaller models can be specifically optimized for retrieval through a systematic study of publicly available training data and data composition strategies. We analyze the effects of training data scale, negative sampling, and data diversity on multilingual retrieval performance. Our results show that increasing data scale yields diminishing returns, whereas incorporating hard negatives is critical for consistent performance improvements. We further find that task diversity in the training data has a greater impact on retrieval quality than language diversity alone. Guided by these findings, we develop a compact multilingual model with approximately 300M parameters that achieves retrieval performance comparable to, and in some cases surpassing, strong 7B-parameter models.
Retrofitting Small Multilingual Models for Retrieval: Matching 7B Performance with 300M Parameters
Lifu Tu, Yingbo Zhou, Semih Yavuz Salesforce AI Research
1 Introduction
Text embedding models (Devlin et al., 2019; Xue et al., 2021; Izacard et al., 2022) encode semantic meaning into dense vectors, enabling efficient retrieval of relevant documents from large corpora. Their importance has increased with the rise of retrieval-augmented generation (RAG) systems (Gao et al., 2024), where retrieval quality directly affects downstream performance. Recent benchmarks such as MMTEB (Enevoldsen et al., 2025), which evaluates over 100 tasks across more than 250 languages, show that large compact multilingual models can perform well. However, for retrieval—the most common use case—small models still trail billion-parameter LLMs on standard retrieval metrics.
In this work, we aim to significantly enhance the retrieval performance of small-scale multilingual embedding models. To accomplish this, we investigate a strategy for retrofitting compact models.Our approach involves systematically exploring several key factors that influence the effectiveness of multilingual embedding training, including the choice of data sources, the scale of training data, and negative sampling strategies. Through extensive experiments, we uncover several important findings. First, we demonstrate that carefully curated synthetic multilingual data, even at small scale (a few thousand examples), yields substantial improvements in retrieval performance. Second, we observe that increasing the amount of training data initially leads to substantial performance gains, but improvements diminish beyond a certain scale, indicating saturation. Third, mining hard negatives with a strong backbone embedding model consistently boosts retrieval accuracy across languages. Finally, our analysis highlights the importance of both task and language diversity in training data, with task diversity having a particularly strong impact on downstream performance.
By prioritizing training strategy over model size, our 300M-parameter model matches or surpasses the retrieval performance of 7B-parameter models on the MMTEB multilingual benchmark, achieving a score of 60.56. Through a systematic study of publicly available data resources and data composition  [^1] In contrast, for concurrent strong models such as EmbeddingGemma (Vera et al., 2025) and Qwen3-Embedding (Zhang et al., 2025), the training data sources and data composition effects are not clearly documented., we show that the 3.5-point improvement over the baseline is driven by three key factors: (i) the use of synthetic multilingual query–document pairs generated from the public mC4 corpus (Xue et al., 2021) with GPT-4o-mini; (ii) the adoption of hard negative mining, which proves more effective than naive data scaling; and (iii) a detailed analysis of data diversity, demonstrating that task diversity has a stronger impact on retrieval performance than simply increasing language coverage.
2 Related work
Multilingual Embedding Models.
Recent work has explored using large language models (LLMs) like XLM-R (Conneau et al., 2020) and Mistral-7B (Jiang et al., 2023) as embedding models, demonstrating clear advantages of LLM-based initialization (Wang et al., 2023; Meng et al., 2024; Yu et al., 2024; Lee et al., 2025). While smaller models, such as multilingual-e5-large-instruct, perform well on benchmarks like MMTEB (Enevoldsen et al., 2025), their retrieval performance still lags behind larger 7B+ models. This gap limits the adoption of compact models in retrieval-based applications. In this work, we address this limitation by leveraging efficient data.
Synthetic Data Generation.
Recent work (Xiao et al., 2023; Lee et al., 2025) has explored using LLMs to generate synthetic data for improving multilingual embeddings, with additional filtering techniques (Thakur et al., 2024; Lee et al., 2025) to enhance data quality. In our work, we use clean Wikipedia documents from mC4 (Xue et al., 2021) and GPT-4o-mini to generate synthetic query-document pairs. Our analysis shows that the generated data is high quality and significantly boosts model performance.
Small Embedding Models.
Earlier small embedding models (Devlin et al., 2019; Xiao et al., 2023; Wang et al., 2024) often struggle with retrieval and short-context limitations. Despite recent advancements from concurrent models like EmbeddingGemma (Vera et al., 2025) and Qwen3-Embedding (Zhang et al., 2025)  [^2] Data details for the two models are unclear., the impact of data resource and composition remains under-explored. In this work, we conduct a rigorous analysis of data effects—including hard negatives and the diversity of tasks and languages—to optimize retrieval performance in public models.
3 Method
Data.
We use existing high-quality and diverse English datasets from multiple task categories, including MS MARCO (Nguyen et al., 2016), HotpotQA (Yang et al., 2018), Natural Questions (NQ) (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), and AllNLI from SimCSE (Gao et al., 2021). Additional datasets covering other task types—such as classification, semantic textual similarity (STS), reranking, clustering, and pair classification—are described in Appendix C.
To mitigate the lack of multilingual retrieval data, we construct high-quality synthetic datasets. Specifically, we sample lengthy (100-1000) Wikipedia articles from mC4 (Xue et al., 2021) [^3] https://huggingface.co/datasets/allenai/c4, then genera a question given the Wikipedia documents with GPT-4o-mini. The generated question and the sampled article constitute a text pair to the fine-tuning data. We generate synthetic data across 20 languages [^4] Please see Appendix A.1 for the full language list.. See the discussion on quality of synthetic data in the Appendix D.
Training.
Given a batch of query document pairs, we use the contrastive loss (Gao et al., 2021):
$\scriptstyle-\frac{1}{N}\Bigg(\sum_{i=1}^{N}\log\frac{\exp(\frac{\textit{sim}(h_{q_{i}},h_{d^{+}})}{\tau})}{\exp(\frac{\textit{sim}(h_{q_{i}},h_{d^{+}})}{\tau})+\sum_{j=1}^{K}\exp(\frac{\textit{sim}(h_{q_{i}},h_{d_{j}^{-}})}{\tau})}\Bigg)$
where $q_{i}$ is the $i$-th query of the mini batch, $d^{+}$ is the positive document, and $d_{j}^{-}$ is the negative document. $\tau$ is the temperature term, $N$ is the mini batch size, $K$ is the number of negative documents associated with each query. $h_{q}$ and $h_{d}$ are the representations of the given query $q$ and document $d$, sim is the similarity function. Cosine similarity is used in our experiments. We set $\tau=0.02$ in our experiments. More experimental details are shown in the Appendix E.
Evaluation.
We use MMTEB (Enevoldsen et al., 2025) for multilingual evaluation, which includes over 500 tasks across 250+ languages and multiple domains. We also evaluate on French, Japanese, and other European languages.
4 Results
Model #Params Supported Context Length Emb. Dim. MMTEB(Retrieval)
MGTE 305M 8k 768 56.50
bge-m3 568M 8k 4096 55.59
multilingual-e5-large-instruct 560M 514 1024 57.12
snowflake-l 568M 8k 1024 58.36
Our 305M 8k 768 60.56
gte-Qwen2-1.5B-instruct 1.5B 32k 1536 60.78
SFR-Embedding-Mistral 7B 4k 4096 59.44
gte-Qwen2-7B-instruct 7B 32k 3584 60.08
Table 1: Performance of models in the MTEB (Multilingual) retrieval tasks.
As shown in prior work (Enevoldsen et al., 2025), multilingual-e5-large-instruct is the best-performing publicly available model on the MTEB (Multilingual) benchmark. However, it only supports a context length of approximately 500 tokens. Despite its strong performance, its retrieval score is over 3 points lower than that of gte-Qwen2-1.5B-instruct. Furthermore, as shown in Table 1, on MTEB (Multilingual) retrieval tasks, the best-performing small multilingual model still lags behind larger models by about 2 points. In this work, we present a compact embedding model (about 300M parameters) that achieves the best performance among small multilingual models on retrieval benchmarks—comparable to or even exceeding that of significantly larger models.
5 Analysis
5.1 Data Resource
In this section, we first examine the usefulness of multilingual synthetic data for retrieval. To explore this, we sample 2,000 examples from three different sources: high-quality English retrieval data from MS MARCO (Nguyen et al., 2016), English–French parallel data from OPUS (Tiedemann, 2012), and our own sampled French synthetic query-document pairs. We evaluate using two backbone embedding models: a multilingual model (MGTE (Zhang et al., 2024)) and an English-only model (BGE-en-1.5 (Xiao et al., 2023)). As shown in Table 2, our synthetic data significantly improves multilingual retrieval performance. On the other hand, parallel data contributes little.
We also explore synthetic data generated by other strong publicly multilingual LLMs, such as Qwen/Qwen3-30B-A3B-Instruct-2507  [^5] See Team (2025). While similar trends are observed, improvements for non-English languages are smaller than those obtained with GPT-4o-mini. Therefore, in the following experiments, we use GPT-4o-mini exclusively as the generator for multilingual synthetic data.
Setting all English Multilingual
MGTE 56.50 49.80 67.44
+English Data 58.6 52.51 68.21
+parallel data 57.24 50.68 67.54
+synthetic data 58.83 52.6 68.81
(a) Multilingual Embedding Model
Setting all English Multilingual
BGE-en-1.5 38.02 45.73 25.91
+English Data 39.30 46.26 26.87
+parallel data 37.67 45.76 23.73
+synthetic data 42.53 49.26 31.36
(b) Non-Multilingual Embedding Model
Table 2: Performance on MTEB (Multilingual) when fine-tuning with different data sources. The column ‘all’, ‘English’, and ‘Multilingual’ refer to the average performance across all retrieval tasks, English-only tasks, and tasks involving non-English languages, respectively.
5.2 Data Scale Effect
Building on the observed gains from synthetic data, we examine the effect of data scale by sampling French synthetic retrieval data at sizes from 1k to 16k. As shown in Figure 1  [^6] Results of the mean over five runs., performance improves with more data but exhibits diminishing returns beyond 4k examples  [^7] We hypothesize that this is due to the limited domain coverage of the synthetic retrieval data, which is generated from Wikipedia articles in mC4.. Accordingly, we use 4k synthetic samples per language in subsequent experiments.
Figure 1: Synthetic Data scale experiments.
5.3 Hard Negatives
Recent studies (Lee et al., 2024, 2025) have demonstrated the effectiveness of incorporating hard negatives in training embedding models. In this section, we compare two strategies for selecting hard negatives: (a) vanilla: neegatives are randomly sampled within each mini-batch. (b) hard negatives: negatives are mined using the backbone multilingual embedding model.
As shown in Figure 2  [^8] Results of the mean over five runs., the hard negative strategy consistently improves multilingual retrieval performance. Based on this observation, we adopt the hard negative mining approach for generating synthetic training samples in our experiments.
Figure 2: Hard negatives experiments.
5.4 Task Diversity and Language Diversity
Mul. Eur. Fr Ja
Our (all) 60.56 59.76 68.28 72.29
En-Diverse 59.60 58.22 67.56 72.26
Mul-Syn 59.36 58.61 67.19 71.46
En-Syn + En-Diverse 60.38 59.56 68.05 72.29
Table 3: Average retrieval performance across varying task and language diversity. "Mul.", "Eur.", "Fr", and "Ja" denote average scores on MTEB Multilingual, MTEB Europe, French (9 tasks), and Japanese (10 tasks) retrieval benchmarks, respectively
As shown in Table 3, training our model with all synthetic multilingual data (Mul-Syn) improves MTEB (Multilingual) retrieval performance to 59.36, a substantial gain over the backbone baseline (56.5). To further isolate the effect of data composition, we train a model using only English data but spanning diverse task types (e.g., classification, STS, clustering; En-Diverse). Despite being monolingual, this model performs strongly, underscoring the importance of task diversity.
Combining multilingual synthetic data with English task-diverse data yields our best model, achieving 60.56. Notably, when training with only English synthetic data and English task-diverse data—without any multilingual synthetic data (Syn-En + En-Diverse)—performance drops by only 0.2 points on MTEB (Multilingual) retrieval. To better contextualize this difference, we additionally report results on MMTEB (Europe) and on 9 French and 10 Japanese retrieval tasks (Appendix F). Across these languages, performance differences remain minimal, indicating that task diversity contributes more to retrieval gains than language diversity when retrofitting multilingual embedding models.
5.5 Performance Across Other Task Categories
Table 4 shows that our models achieve over 1 point higher overall performance and outperform the baseline across most task types. This improvement suggest the effectiveness of our training strategy—particularly the incorporation of diverse task types and synthetic multilingual data—in producing more robust and generalizable multilingual embeddings. The consistent gains across multiple task categories suggest that the model is not overfitting to any single task or language, but is instead learning semantically rich representations that transfer well to a variety of scenarios.
Our MGTE
Mean (Type) 52.66 51.51
- Bitext Mining 71.61 71.79
- Classification 58.98 57.17
- Clustering 46.11 44.96
- Inst. Retrieval -0.85 -0.74
- Multilabel Class. 20.84 19.82
- Pair Class. 81.01 80.49
- Reranking 61.70 60.72
- Retrieval 60.56 56.50
- STS 73.96 72.85
Table 4: Comparison with the backbone model (MGTE) on the MTEB (Multilingual) benchmark.
6 Conclusion
In this work, we develop a compact multilingual embedding model (approximately 300M parameters) that achieves retrieval performance comparable to—or even surpassing—that of current strong 7B models. This demonstrates that small models can be effectively used for retrieval, the most critical and practical application of embedding models.
Limitations
In this work, we explore methods to improve the retrieval performance of small multilingual models and demonstrate consistent gains across most task types. However, several limitations remain:
Benchmark
MTEB (Multilingual) covers over 250 languages; however, low-resource languages are primarily represented in task categories such as bitext mining and classification. The evaluation of retrieval performance for these languages remains limited, which constrains our ability to assess the effectiveness of the proposed methods in low-resource settings. This limitation may also explain why the improvement from multilingual synthetic data—compared to English-only synthetic data—is relatively small, as shown in Table 3. Because of the limitation of the existing benchmark, it is hard to evluate performance on very low-resource languages  [^9] Although the used benchmark covers over 250 languages, retrieval performance on low-resource languages is already high, leaving limited room for further improvement..
Efficiency
In this work, we show that multilingual synthetic data effectively improves retrieval performance, but its benefits eventually plateau, indicating a limitation when relying on retrieval data on few domains alone. We further find that incorporating diverse non-retrieval task data leads to additional gains. Scaling the joint use of diverse non-retrieval task data and multilingual synthetic retrieval data is a promising direction for future research.
Appendix A Details on Synthetic Data Generation
A.1 Languages for Synthetic Data
List of language ISO-639 code (https://cloud.google.com/translate/docs/languages): en, fr, es de,zh, it, ja, ko, fa, hi, id, ar, bn, fi, sw, te, th, jv, ms, sq.
A.2 Template for Synthetic Data Generation
. Give a sampled Wikipedia documents {d} and target language {lang}, the following template is used to generate a query:
You are a curious AI assistant, please generate one specific and valuable question based on the following text. If the text is not {lang}, please reply with Non-{lang}. The generated question should revolve around the core content of this text, and avoid using pronouns (e.g., ’this’). Note that you should generate only one question, without including additional content:
{d}
Appendix B Models
Below is a list of models discussed in this work:
  • MGTE (Zhang et al., 2024) [^10] Alibaba-NLP/gte-multilingual-base
  • BGE-en-1.5 (Xiao et al., 2023) [^11] BAAI/bge-large-en-v1.5
  • bge-m3 (Chen et al., 2024) [^12] BAAI/bge-m3
  • multilingual-e5-large-instruct (Wang et al., 2024)131313intfloat/multilingual-e5-large-instruct
  • snowflake-l (Yu et al., 2024) [^14] Snowflake/snowflake-arctic-embed-l-v2.0
  • gte-Qwen2-1.5B-instruct (Li et al., 2023) [^15] Alibaba-NLP/gte-Qwen2-1.5B-instruct
  • SFR-Embedding-Mistral (Meng et al., 2024) [^16] Salesforce/SFR-Embedding-Mistral
  • gte-Qwen2-7B-instruct (Li et al., 2023) [^17] Alibaba-NLP/gte-Qwen2-7B-instruct
In this work, the AI assistant ChatGPT is used solely for writing and language polishing.
Appendix C Additional Data Details
The following are additional English data samples from the training split of MTEB, excluding data used in the official MTEB evaluation.
  • Classification: AmazonPolarity, AmazonReviews, Imdb, MTOPDomainClassification,AmazonReviewsClassification
  • STS: STS12, STS22, STSBenchmark
  • Reranking: SciDocsRR and StackOverflowDupQuestions
  • Clustering: arXIV, bioRxiv, medRxiv
  • Pair Classification: SprintDuplicateQuestions, TwitterURLCorpus
Appendix D Quality of synthetic data
In order to avoid "model collapse" when training on synthetic data, we take several ways to improve the quality of the synthetic data:
  • High-quality multilingual Wikipedia documents (e.g., mC4; Xue et al., 2021) are used as clean source material for synthetic data generation.
  • For each ⟨question, document⟩ pair, only the question is generated by a large language model, while the document is drawn from existing sources. We further restrict the document length to between 100 and 1,000 tokens.
  • To assess question quality, we randomly sampled 50 questions for each language and manually evaluated them. Surprisingly, most questions are well-formed and relevant to the given documents.
Appendix E Implementation Details
We use a batch size of 1024 and a learning rate of 1e-5, with 100 warm-up steps followed by linear decay. The maximum sequence lengths are set to 256 for queries and 512 for documents. LoRA adapters with a rank of $r$ = 8 are applied. Each query-document pair is trained with 7 hard negatives, all drawn from the same task.
Appendix F Additional Retrieval Tasks
9 French retrieval tasks:
French subset of the following 9 tasks: AlloprofRetrieval, BSARDRetrieval, FQuADRetrieval, SyntecRetrieval, BelebeleRetrieval, MIRACLRetrievalHardNegatives, PublicHealthQA, StatcanDialogueDatasetRetrieval, WebFAQRetrieval
10 Japanese retrieval tasks:
Japanese subset of the following 10 tasks: JaGovFaqsRetrieval, JaqketRetrieval, JaQuADRetrieval, NLPJournalAbsIntroRetrieval, NLPJournalTitleAbsRetrieval, NLPJournalTitleIntroRetrieval, BelebeleRetrieval, MIRACLRetrievalHardNegatives, MrTidyRetrieval, WebFAQRetrieval.